Status Report

NOAA GOES SATOPS Morning Report 22 Jan 2001

By SpaceRef Editor
January 22, 2001
Filed under ,

SATOPS MORNING REPORT January 22, 2001

POES

NOAA-16 rev 1697 / W at 1757Z on January 19: Commanded VHF Transmitter
2 (VTX2) off.

NOAA-16 rev 1698 / W at 1939Z on January 19: Commanded VHF Transmitter
1 (VTX1) on.

NOAA-14 rev 31219 / W at 2053Z on January 19: Commanded VHF Transmitter
2 (VTX2) off.

NOAA-16 rev 1718 / W at 0612z on January 21: A ground system timing
problem caused the loss of three data sets, which had been received without
incident on the IMUX ground recroder during this pass. G2B, G3B and G5B
were all affected by this problem.  G3B and G5B were recovered on
rev 1719 at 0752z. G2B is not recoverable due to spacecraft recorder scheduling
constraints.

NOAA-16 rev 01726 / W at 1919Z on January 21: DUAL LINK APT test completed. 
VTX1 AGC levels averaged -98 dbm, so the XMTR was not powered off. 
CDA personnel indicated that this signal should be usable for average receiver
capabilities.

 

DMSP

OPS49 rev 05714 / POGO at 2031z on January 19:  Nominal commanding
accomplished to put spacecraft back into single processor operating mode.

At 1404z on January 20: DMSP S20 Launch scrubbed and delayed for 24
hours: Unable to command Signal Conditioning Unit (SCU) on spacecraft from
Payload Integration Test Facility.

Launch Delay: At 1408z on January 21 the DMSP S20 Launch, scheduled
for 1358Z, was scrubbed at the end of the launch window. A second stage
oxidizer pre-valve status halted the countdown at T-19 sec. Count was restarted
once the valve status cleared, but there was insufficient time to restart
and execute countdown sequence before the end of the launch window. 
Officials have decided to further postpone the Titan 2 launch and will
forego making a launch attempt on Monday morning due to detection and investigation
of a fuel leak on the launch pad. Liftoff is now scheduled for Tuesday
at 1358 GMT (8:58 a.m. EST; 5:58 a.m. PST).

 

GOES

No change in the status of the GOES spacecraft.

1833Z on January 20: GOES-10 image frame break due to east/west address
error.

0104Z on January 21: GOES-10 image frame break due to loss of signal
strength.

1144Z on January 21: GOES-8 image frame break due to high bit error
rate.

0220Z on January 22: GOES-10 image frame break due to loss of signal
strength.

GOES operations were otherwise nominal over the past 72 hours.

 

Questions or comments to [email protected]

SpaceRef staff editor.